OpenStudy (matt6288):

Use the quadratic formula to solve the quadratic equation x2 − 8x − 4 = 0

OpenStudy (anonymous):

Is that x squared? Like, x^2?

OpenStudy (matt6288):

\[4x 2\sqrt{5}\] its somthing like that?..

OpenStudy (anonymous):

Cool, thanks. :D I'm doing it right now, haha.

OpenStudy (matt6288):

its no fun haha

OpenStudy (anonymous):

Awh it's easy once you get used to it. :D

OpenStudy (matt6288):

then whats the answer lmfaoo

OpenStudy (accessdenied):

Quadratic formula is pretty much a "get the formula and your equation and then plug+play" sort of thing. Once you have the equation, \(ax^2 + bx + c = 0\), the solutions are given by the equation: \[ x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \]